# MIRI spectrophotometry of GN-z11: Detection and nature of an optical red continuum component

### v1

@@ 3.3 SED fitting
In order to investigate the nature of GN-z11, we have analysed its SED based on the available spectro-photometric data. In particular, the SED constructed with the NIRCam+MIRI photometry and the NIRSpec+MRS spectroscopy was fitted with two codes, CIGALE v2025.0 (Burgarella2005; Noll2009; Boquien2019) and SYNTHESIZER-AGN (Perez-Gonzalez+03_Synthesizer; Perez-Gonzalez+08_Synthesizer). Since the F1000W flux shows a lower significance compared to the other MIRI values, we exclude F1000W (and the upper limit for F2100W) from the SED fitting to prevent biasing the results. However, we will extract the modelled fluxes at these filters to evaluate their consistency.
@@
The SED fitting from CIGALE makes use of the UV+optical emission lines and the MIRI fluxes. In addition, as the current version of CIGALE does not use spectroscopic data, we have constructed pseudo-continuum fluxes by integrating the re-normalised NIRSpec/Prism spectrum in 750 bins. The emission lines were assigned separate bins to avoid contaminating the pseudo-continuum values. These UV pseudo-continuum, which isolate the effect of the emission lines and trace the galaxy continuum emission, are considered during the SED fitting along with the MIRI fluxes. In our CIGALE model, the stellar component is fitted using the stellar population templates from the Binary Population and Spectral Synthesis library (BPASSv2.2, Stanway+18) with = 0.2 (following Alvarez-marquez+25) and a Chabrier initial mass function (IMF, Chabrier+03) with an upper stellar mass limit of 100 . Nebular emission is included via Cloudy models (Ferland+17), assuming = 0.2 , an electron density of 1000 cm-3, and an ionisation parameter of = (following Bunker+23). Dust extinction is assumed to follow the attenuation curve from Calzetti+00 with 0.3 mag, constrained by the Balmer line ratios derived from previous NIRSpec and MRS analysis (Bunker+23; Alvarez-marquez+25). We have also considered the potential presence of a non-stellar component (i.e., type 1 or 2 AGN) by adding an AGN component modelled as described by Fritz+06. This model assumes a broken power-law point source as the ionization source and the presence of a smooth dusty torus. Since our dataset covers up to rest-frame 0.86 m, we cannot fully constrain the full model parameter space. Therefore, for simplicity, in this work we fixed an opening angle of 60°and a / = 10, following the results found for high- luminous and massive QSOs (Bosman+25) while we allow the radial and angular indices range between = [-1, -0.5] and = [0, 6], respectively. We tested both type 1 and type 2 scenarios by forcing the angle of the line of sight to the AGN polar axis be ¡20° and ¿70°, respectively.
@@
For the SED fitting carried out with SYNTHESIZER-AGN, we used the re-normalised NIRSpec/Prism spectrum, the MRS optical lines and the MIRI fluxes. Throughout the fitting, we consider the stellar population models described by Bruzual&Charlot+03, hereafter BC03, assuming a Chabrier+03 IMF with stellar masses between 0.1 and 100 . The nebular continuum and emission lines are modelled using Cloudy c23.01 (Gunasekera+23) predictions for a variety of ionizing SEDs (parametrised by a range of effective temperatures (8000 120000 K) and ionizing photon flux ( 1060 s-1)), and a variety of electron densities (log(/cm-3) = 2 4) and metallicities (0.02 0.2 , with solar abundances for individual elements). The dust extinction is described by the Calzetti+00 attenuation law where we constrained the to be lower than 0.30 mag, as for the CIGALE fitting. The AGN emission is included following the QSO composite spectrum created by combining the UV+optical and near-IR templates from Selsing+16 (up to rest- 1.13 m) and Glikman+06 (rest- 1.13 3.5 m), respectively. This empirical template is added to the fitting procedure to represent average accretion disk and dust emission in luminous type 1 AGNs. Additionally, we also consider the presence, without energy balance, of an additional dusty torus to reproduce the observed red SED shape. This component is defined by using self-consistent AGN torus model from Polletta+06; Polletta+07, which is a highly obscured ( = 4) QSO with UV emission coming from scattered light and IR emission coming from hot/warm dust.
@@
To quantify the dependence of the physical parameters when using different stellar population models, we compared the CIGALE results considering both the BC03 and BPASSv2.2 stellar templates. The BPASSv2.2 models predicts stronger UV emission due to binary evolution, allowing stellar populations older than 10 Myr to still produce significant ionizing flux. This yields typically older ages, higher SFRs, and lower stellar masses, while BC03 favours younger and more massive solutions for GN-z11. However, these differences are 10 15 and, in all the cases, smaller than those introduced by assuming different star-formation histories (SFHs). We have also examined the potential effect of adopting a different extinction law during the SED fitting. For this comparison, we have tested the Charlot&Fall2000 attenuation curve, in which the young stars and the nebular emission suffer an additional attenuation compared with the older stars ( 10 Myr). We find that, in our case, the SED fitting using Charlot&Fall2000 lead to stellar masses only 0.1 dex smaller than those assuming Calzetti+00 while variations in other physical parameters (e.g., SFR, ) are lower than 10. We therefore argue that neither the choice of the stellar population templates or the attenuation curve affects any conclusion drawn from our analysis.
@@ 4.1 Continuum-only fluxes and colour-colour diagrams
The large spectro-photometric dataset available for GN-z11 combining NIRCam, NIRSpec, MIRI and MRS allows us to trace both the nebular lines and the continuum from rest- UV up to 0.86 m. By removing the nebular line contributions from the aperture photometry, we obtain continuum-only fluxes, which allow us to trace the GN-z11 underlying continuum emission (see Sect. 3.2).
@@
The continuum-only results show (see Table 1) how, although the MIRI/F560W value agree within the uncertainties with the continuum level expected from previous SED fittings based on UV-only data (Jiang+21; Tacchella+23_GNz11), our F770W photometric value exceeds their predicted flux. In particular, once we take into account the difference with the photometry from Tacchella+23_GNz11 (see App. A), our F770W continuum level (212 20 nJy) is 70 larger (5) than the SED fitting expected value when using UV-only photometry ( 120 nJy). This excess at 0.66 m rest-frame could be hinting at the presence of a non-negligible old stellar population or a red component that would be strongly emitting at optical-red and near-IR wavelengths.
@@
A first step to investigate the nature of GN-z11, and this possible red excess, is to use classic colour-colour diagrams covering from rest-frame UV (0.17 m) to optical-red (0.66 0.86 m) wavelengths. Figure 2 displays the diagrams for the UV (F200W F444W), optical (F444W F560W) and optical-red (F560W F770W and F770W F1000W) colours derived from the continuum-only fluxes. In this figure we compare these continuum-only colours to those expected from stellar populations models and those from intrinsically red objects (e.g., QSO, AGNs and LRDs). In particular, we have considered the stellar populations templates from BPASSv2.2 for continuum and instantaneous SFHs and a range of stellar ages limited to the age of the Universe at = 10.6 (i.e., 440 Myr). Their associated nebular continua are also included using the corresponding Cloudy models for those ages and SFHs (Ferland+17). For completeness, we displayed the age evolution tracks for metallicities 0.05 and 0.2 . Green diamonds represent the colours of the nuclear ultraviolet-optical spectrum of Mrk 3. This low- galaxy is classified as a heavily obscured Seyfert 2 nucleus (Guainazzi+16), with a hidden BLR identified by the presence of broad Balmer lines in polarised light (Miller1990). Its spectrum, obtained with the HST/STIS spectrograph using an aperture size of 0.2″(i.e., 50 pc at the distance of Mrk3, Spinelli+06), represents the SED of a luminous type 2 AGN while minimizing the potential contamination due to the stellar light that could be affecting larger apertures. Moreover, its line ratios (O3H = , Ne3O2 = and O32 = 0.9), agree with those found in high- AGNs (Mazzolari+24; Mazzolari+25). Orange diamonds correspond to the colours measured for the template of luminous blue QSOs at intermediate redshifts (1 ¡ ¡ 2) constructed from the simultaneous ultraviolet to near-infrared spectroscopy (Glikman+06; Selsing+16). Purple diamonds represent the colours derived for the NIRSpec+MRS spectrum of the = 7.1 QSO J1120+0641, whose BLR and dust emission has been found to be consistent with those of lower-redshift quasars (Bosman2024; Bosman+25). We have filled the gap between the NIRSpec and MRS data (i.e., 3.12 4.9 m) following its best-fit SED shape ( ; Bosman+25). Brown diamonds show the colour derived from the spectra of the singular little red dot (LRD) presented by deGraaff+25, known as ‘The Cliff’. This object presents the strongest Balmer break among LRDs and shows H absorption, which are explained under the ‘black hole star’ (BH*) scenario in which dense gas surrounds a powerful ionising source (see also Naidu+25 and Taylor+25). Finally, the red and blue diamonds display the colour derived for an updated stacked spectra from PerezGonzalez-LRD2024, combining 150 nearby and high- LRDs, and the COSMOS-Webb median-stacked LRD from Akins+25, respectively. The spectra of all these sources together with the GN-z11 JWST photometry are presented in Figure 3.
@@
The continuum-only fluxes show an SED blue in the rest-UV (F200W F444W = 0.30 0.05 mag) with a flattening in up to rest-frame 0.5 m (F444W F560W = 0.12 0.21 mag). Beyond [O III] 5008, the SED appears to rise, with colours F560W F770W = 0.33 0.24 mag and F770W F1000W = 0.18 0.40 mag (0.26 0.24 and 0.33 0.40 mag, respectively, if we consider the impact of the potential broad hydrogen Balmer lines components, see Sect. 3.2). Although the rest- UV and optical colours can be reproduced by a young ( 10 Myr) stellar burst, it fails to explain the optical-red colours (e.g., F560W F770W and F770W F1000W, see Figure 2). On the other hand, the Seyfert 2 spectrum can produce the optical-red continuum-only colours observed with MIRI but fails to explain the UV-blue colour F200W F444W. Finally, the LRD stacked and ‘The Cliff’ spectra are too red, specially beyond 0.4 m, and they are only marginally consistent with F770W F1000W due to the large uncertainty in the F1000W flux. These results, based on the traditional colour-colour diagrams, suggest that while the UV emission from GN-z11 can be produced by a young star-formation, the presence of an extra red source is needed to explain its rest-frame optical-red colours. Note however that in these colour-colour diagrams, the stellar plus nebular continuum models do not contain any contribution due to small amounts of hot dust emitting at optical-red wavelengths. Likewise, mixed stellar populations, that would produce optical colours different from those of (single age) instantaneous bursts are not represented.
@@
In the following sections we perform a more detailed modelling using different SED fitting codes with different scenarios (e.g., SFHs, AGN; see Sect. 3.3) to investigate the origin behind these UV, optical and optical-red colours. Besides, SED fitting will allow us to model both the continuum and the emission lines simultaneously, providing a more complete view of the physical mechanisms driving the emission in GN-z11.
@@ 4.2 SF-only SED modelling
As a first approach for the CIGALE SED fitting, we considered that GN-z11 can be modelled by a stellar population following a tau-decaying SFH with varying in the [1, 5000 Myr] range, and with ages up to 440 Myr (i.e., age of Universe at = 10.603). Figure 4 shows the best-fit SED obtained from these adopted model parameters and the comparison with the observed NIRCam+MIRI fluxes. The Bayesian analysis results in an almost constant SFH ( = 2500 1500 Myr) and a stellar population of = (5.6 0.6) 108 and an age of = 21 3 Myr. Although for this model, we obtain H and [O III] 5008 that are in agreement with those measured with the MRS, the modelled F770W flux (i.e., 176 1 nJy) is underestimated by 45, 5 below the value observed with MIRI (see Table 1). The blue optical continuum of this young stellar population makes this difference more pronounced in F1000W, modelling a flux 44 fainter (2 considering the large F1000W uncertainties). For completeness, we have tested an exponentially growing SFH (negative ) obtaining consistent results.
@@
Since a single young stellar population is not able to justify the UV and optical continuum and the nebular emission of this galaxy, we consider the possibility of having a mix of different stellar populations. In recent years, a number of works have proposed that stellar populations of different ages are required to fit the UV to optical SED of high- galaxies. Tacchella+23 found the presence of old stellar populations in ¿7 galaxies using SED fitting based on NIRSpec+NIRCam data. Underlying 100 Myr stellar populations have been also found in five 5¡¡9 galaxies in which the SEDs are dominated by blue emission produced in young bursty star-forming regions (Gimenez-Arteaga+23). Some authors have found that a double power-law describes better the SFHs in high- galaxies, as single population models can be out-shone by young UV-bright stars, yielding to lower stellar masses (Gimenez-Arteaga+24) . In addition, stellar populations older than 10 Myr produce a redder optical continuum, in accordance with the MIRI colours found in GN-z11 (see Fig. 2).
@@
Therefore, we have also considered a mixed stellar component consisting of a young population ( 10 Myr) with a constant star formation and a more mature one (20 440 Myr) formed in a past burst. In this scenario, GN-z11 can be interpreted as a young compact burst in a mature host galaxy. Figure 4 displays the best-fit SED obtained under this approach and the residuals with the observed NIRCam+MIRI fluxes. Under this mixed stellar population scenario, GN-z11 is composed of a young (4 1 Myr) stellar population that represents the 12 of the total stellar mass (1.5 0.1 109 ), while the more mature one has an average age of 63 23 Myr. These values are consistent with the double-component SED fitting carried out in Tacchella+23_GNz11, where the authors found a point-source with a young stellar population ( 11 Myr) along with a more massive and mature one ( 109 and 35 Myr) dominating the extended emission. Although the near-IR emission produced by the mature stellar population increases the modelled flux in F770W up to 202 5 nJy, it is still 3 fainter than the measured value (see Table 2). Even if we add the contribution of the upper limit of the potential broad component of H (2 3 1018 erg s-1 cm-2, Alvarez-marquez+25), the flux in F770W would increase only by 18 27 nJy and would be still incompatible with the MIRI data by almost 2 (see Table 2). Despite the increase in the modelled flux in F1000W (i.e., 130 6 nJy) driven by the optical and near-IR emission from the older stellar population, we are not recovering the observed MIRI flux (i.e., 249 77 nJy).
@@
Although assuming either a single or mixed stellar population provides a good SED fit up to rest-frame 0.5 m, both approaches fail to reproduce the F770W and F1000W photometry. As discussed in Sect. 4.1, the significant increase in the SED of GN-z11 beyond 0.66 m cannot be reproduced solely by a combination of different stellar populations. This result suggests the presence of an additional, intrinsically red, component in GN-z11.

### v2

@@ 3.3 SED fitting
In order to investigate the nature of GN-z11, we analysed its SED based on the available spectro-photometric data. In particular, the SED constructed with the NIRCam+MIRI photometry and the NIRSpec+MRS spectroscopy was fitted with two codes, CIGALE v2025.0 (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019) and SYNTHESIZER-AGN (Pérez-González et al., 2003, 2008). Since the F1000W flux shows a lower significance compared to the other MIRI values, we excluded F1000W (and the upper limit for F2100W) from the SED fitting to prevent biasing the results. However, we extracted the modelled fluxes at these filters to evaluate their consistency.
@@
The SED fitting from CIGALE makes use of the UV+optical emission lines and the MIRI fluxes. In addition, as the current version of CIGALE does not use spectroscopic data, we constructed pseudo-continuum fluxes by integrating the re-normalised NIRSpec/Prism spectrum in 750 bins. The emission lines were assigned separate bins to avoid contaminating the pseudo-continuum values. These UV pseudo-continuums, which isolate the effect of the emission lines and trace the galaxy continuum emission, were considered during the SED fitting along with the MIRI fluxes. In our CIGALE model, the stellar component was fitted using the stellar population templates from the Binary Population and Spectral Synthesis library (BPASSv2.2, Stanway and Eldridge 2018) with = 0.2 (following Álvarez-Márquez et al. 2025) and a Chabrier IMF (Chabrier, 2003) with an upper stellar mass limit of 100 . Nebular emission was included via Cloudy models (Ferland et al., 2017), assuming = 0.2 , an electron density of 1000 cm-3, and an ionisation parameter of = (following Bunker et al. 2023). Dust extinction was assumed to follow the attenuation curve from Calzetti et al. (2000) with 0.3 mag, constrained by the Balmer line ratios derived from previous NIRSpec and MRS analysis (Bunker et al., 2023; Álvarez-Márquez et al., 2025). We also considered the potential presence of a non-stellar component (i.e. type 1 or 2 AGN) by adding an AGN component modelled as described by Fritz et al. (2006). This model assumes a broken power-law point source as the ionisation source and the presence of a smooth dusty torus. Since our dataset covers up to rest-frame 0.86 m, we cannot fully constrain the full model parameter space. Therefore, for simplicity, in this work we fixed an opening angle of 60°and a / = 10, following the results found for high- luminous and massive quasi-stellar objects (QSOs; Bosman et al., 2025), while we allowed the radial and angular indices to range between = [-1, -0.5] and = [0, 6], respectively. We tested both type 1 and type 2 scenarios by forcing the angle of the line of sight to the AGN polar axis to be ¡20° and ¿70°, respectively.
@@
For the SED fitting carried out with SYNTHESIZER-AGN, we used the re-normalised NIRSpec/Prism spectrum, the MRS optical lines, and the MIRI fluxes. Throughout the fitting, we considered the stellar population models described by Bruzual and Charlot (2003), hereafter BC03, assuming a Chabrier (2003) IMF with stellar masses between 0.1 and 100 . The nebular continuum and emission lines were modelled using Cloudy c23.01 (Gunasekera et al., 2023) predictions for a variety of ionising SEDs (parametrised by a range of effective temperatures (8000 120000 K) and ionising photon flux ( 1060 s-1)), and a variety of electron densities (log(/cm-3) = 2 4) and metallicities (0.02 0.2 , with solar abundances for individual elements). The dust extinction is described by the Calzetti et al. (2000) attenuation law whereby we constrained the to be lower than 0.30 mag, as for the CIGALE fitting. The AGN emission was included following the QSO composite spectrum created by combining the UV+optical and near-IR templates from Selsing et al. (2016) (up to rest- 1.13 m) and Glikman et al. (2006) (rest- 1.13 3.5 m), respectively. This empirical template was added to the fitting procedure to represent the average accretion disc and dust emission in luminous type 1 AGNs. Additionally, we also considered the presence, without an energy balance, of an additional dusty torus to reproduce the observed red SED shape. This component was defined by using self-consistent AGN torus model from Polletta et al. (2006, 2007), which is a highly obscured ( = 4) QSO with UV emission coming from scattered light and IR emission coming from hot and/or warm dust.
@@
To quantify the dependence of the physical parameters when using different stellar population models, we compared the CIGALE results considering both the BC03 and BPASSv2.2 stellar templates. The BPASSv2.2 models predict stronger UV emission due to binary evolution, allowing stellar populations older than 10 Myr to still produce significant ionising flux. This typically yields older ages, higher SFRs, and lower stellar masses, while BC03 favours younger and more massive solutions for GN-z11. However, these differences are 10 15 and, in all cases, smaller than those introduced by assuming different star formation histories (SFHs). We also examined the potential effect of adopting a different extinction law during the SED fitting. For this comparison, we tested the Charlot and Fall (2000) attenuation curve, in which the young stars and the nebular emission suffer an additional attenuation compared with the older stars ( 10 Myr). We find that, in our case, the SED fitting using Charlot and Fall (2000) lead to stellar masses only 0.1 dex lower than those assuming Calzetti et al. (2000), while variations in other physical parameters (e.g. SFR, ) are lower than 10. We therefore argue that neither the choice of the stellar population templates nor that of the attenuation curve affects any conclusion drawn from our analysis.
@@ 4.1 Continuum-only fluxes and colour-colour diagrams
The large spectro-photometric dataset available for GN-z11 combining NIRCam, NIRSpec, MIRI, and MRS allows us to trace both the nebular lines and the continuum from rest- UV up to 0.86 m. By removing the nebular line contributions from the aperture photometry, we obtained continuum-only fluxes, which allowed us to trace the GN-z11 underlying continuum emission (see Sect. 3.2).
@@
The continuum-only results show (see Table 1) how, although the MIRI/F560W value agree within the uncertainties with the continuum level expected from previous SED fittings based on UV-only data (Jiang et al., 2021; Tacchella et al., 2023a), our F770W photometric value exceeds their predicted flux. In particular, once we take into account the difference with the photometry from Tacchella et al. (2023a) (see App. A), our F770W continuum level (212 20 nJy) is 70 larger (5) than the SED fitting expected value when using UV-only photometry ( 120 nJy). This excess at 0.66 m rest-frame could be hinting at the presence of a non-negligible old stellar population or a red component that would be strongly emitting at optical-red and near-IR wavelengths.
@@
A first step to investigate the nature of GN-z11, and this possible red excess, is to use classic colour-colour diagrams covering from rest-frame UV (0.17 m) to optical-red (0.66 0.86 m) wavelengths. Figure 2 displays the diagrams for the UV (F200W F444W), optical (F444W F560W), and optical-red (F560W F770W and F770W F1000W) colours derived from the continuum-only fluxes. In this figure we compare these continuum-only colours to those expected from stellar populations models and those from intrinsically red objects (e.g. QSO, AGNs, and LRDs). In particular, we considered the stellar population templates from BPASSv2.2 for continuum and instantaneous SFHs and a range of stellar ages limited to the age of the Universe at = 10.6 (i.e. 440 Myr). Their associated nebular continua were also included using the corresponding Cloudy models for those ages and SFHs (Ferland et al., 2017). For completeness, we displayed the age evolution tracks for metallicities 0.05 and 0.2 . Green diamonds represent the colours of the nuclear ultraviolet-optical spectrum of Mrk 3. This low- galaxy is classified as a heavily obscured Seyfert 2 nucleus (Guainazzi et al., 2016), with a hidden BLR identified by the presence of broad Balmer lines in polarised light (Miller and Goodrich, 1990). Its spectrum, obtained with the HST/STIS spectrograph using an aperture size of 0.2″(i.e. 50 pc at the distance of Mrk3, Spinelli et al. 2006), represents the SED of a luminous type 2 AGN, while minimising the potential contamination due to the stellar light that could be affecting larger apertures. Moreover, its line ratios (O3H = , Ne3O2 = , and O32 = 0.9), agree with those found in high- AGNs (Mazzolari et al., 2024, 2025). Orange diamonds correspond to the colours measured for the template of luminous blue QSOs at intermediate redshifts (1 ¡ ¡ 2) constructed from the simultaneous ultraviolet to near-infrared spectroscopy (Glikman et al., 2006; Selsing et al., 2016). Purple diamonds represent the colours derived for the NIRSpec+MRS spectrum of the = 7.1 QSO J1120+0641, whose BLR and dust emission has been found to be consistent with those of lower-redshift quasars (Bosman et al., 2024, 2025). We filled the gap between the NIRSpec and MRS data (i.e. 3.12 4.9 m) following its best-fit SED shape ( ; Bosman et al. 2025). Brown diamonds show the colour derived from the spectra of the singular little red dot (LRD) presented by de Graaff et al. (2025), known as ‘The Cliff’. This object presents the strongest Balmer break among LRDs and shows H absorption, which are explained under the ‘black hole star’ (BH*) scenario in which dense gas surrounds a powerful ionising source (see also Naidu et al. 2025 and Taylor et al. 2025). Finally, the red and blue diamonds display the colour derived for an updated stacked spectra from Pérez-González et al. (2024b), combining 150 nearby and high- LRDs, and the COSMOS-Webb median-stacked LRD from Akins et al. (2025), respectively. The spectra of all these sources together with the GN-z11 JWST photometry are presented in Figure 3.
@@
The continuum-only fluxes show an SED blue in the rest-UV (F200W F444W = 0.30 0.05 mag) with a flattening in up to rest-frame 0.5 m (F444W F560W = 0.12 0.21 mag). Beyond [O III] 5008, the SED appears to rise, with colours F560W F770W = 0.33 0.24 mag and F770W F1000W = 0.18 0.40 mag (0.26 0.24 and 0.33 0.40 mag, respectively, if we consider the impact of the potential broad hydrogen Balmer lines components, see Sect. 3.2). Although the rest- UV and optical colours can be reproduced by a young ( 10 Myr) stellar burst, it fails to explain the optical-red colours (e.g. F560W F770W and F770W F1000W, see Figure 2). On the other hand, the Seyfert 2 spectrum can produce the optical-red continuum-only colours observed with MIRI but fails to explain the UV-blue colour F200W F444W. Finally, the LRD stacked and The Cliff spectra are too red, especially beyond 0.4 m, and they are only marginally consistent with F770W F1000W due to the large uncertainty in the F1000W flux. These results, based on the traditional colour-colour diagrams, suggest that while the UV emission from GN-z11 can be produced by a young star formation burst, the presence of an extra red source is needed to explain its rest-frame optical-red colours. Note, however, that in these colour-colour diagrams, the stellar plus nebular continuum models do not contain any contribution due to small amounts of hot dust emitting at optical-red wavelengths. Likewise, mixed stellar populations that would produce optical colours different from those of (single age) instantaneous bursts are not represented.
@@
In the following sections, we perform more detailed modelling using different SED fitting codes with different scenarios (e.g. SFHs, AGN; see Sect. 3.3) to investigate the origin of these UV, optical, and optical-red colours. In addition, SED fitting also allows us to model both the continuum and the emission lines simultaneously, providing a more complete view of the physical mechanisms driving the emission in GN-z11.
@@ 4.2 SF-only SED modelling
As a first approach for the CIGALE SED fitting, we considered that GN-z11 can be modelled by a stellar population following a tau-decaying SFH with varying in the [1, 5000 Myr] range, and with ages up to 440 Myr (i.e. age of Universe at = 10.603). Figure 4 shows the best-fit SED obtained from these adopted model parameters and the comparison with the observed NIRCam+MIRI fluxes. The Bayesian analysis results in an almost constant SFH ( = 2500 1500 Myr) and a stellar population of = (5.6 0.6) 108 and an age of = 21 3 Myr. Although for this model, we obtain H and [O III] 5008 that are in agreement with those measured with the MRS, the modelled F770W flux (i.e. 176 1 nJy) is underestimated by 45, 5 below the value observed with MIRI (see Table 1). The blue optical continuum of this young stellar population makes this difference more pronounced in F1000W, modelling a flux 44 fainter (2 considering the large F1000W uncertainties). For completeness, we have tested an exponentially growing SFH (negative ), obtaining consistent results.
@@
Since a single young stellar population is not able to justify the UV and optical continuum and the nebular emission of this galaxy, we considered the possibility of having a mix of different stellar populations. In recent years, a number of works have proposed that stellar populations of different ages are required to fit the UV to optical SED of high- galaxies. Tacchella et al. (2023b) found the presence of old stellar populations in ¿7 galaxies using SED fitting based on NIRSpec+NIRCam data. Underlying 100 Myr stellar populations have also been found in five 5¡¡9 galaxies in which the SEDs are dominated by blue emission produced in young bursty star-forming regions (Giménez-Arteaga et al., 2023). Some authors have found that a double power law better describes the SFHs in high- galaxies, as single population models can be out-shone by young UV-bright stars, yielding to lower stellar masses (Giménez-Arteaga et al., 2024) . In addition, stellar populations older than 10 Myr produce a redder optical continuum, in accordance with the MIRI colours found in GN-z11 (see Fig. 2).
@@
Therefore, we also considered a mixed stellar component consisting of a young population ( 10 Myr) with a constant star formation and a more mature one (20 440 Myr) formed in a past burst. In this scenario, GN-z11 can be interpreted as a young compact burst in a mature host galaxy. Figure 4 displays the best-fit SED obtained under this approach and the residuals with the observed NIRCam+MIRI fluxes. Under this mixed stellar population scenario, GN-z11 is composed of a young (4 1 Myr) stellar population that represents 12 of the total stellar mass (1.5 0.1 109 ), while the more mature one has an average age of 63 23 Myr. These values are consistent with the double-component SED fitting carried out in Tacchella et al. (2023a), in which the authors found a point-source with a young stellar population ( 11 Myr) along with a more massive and mature one ( 109 and 35 Myr) dominating the extended emission. Although the near-IR emission produced by the mature stellar population increases the modelled flux in F770W up to 202 5 nJy, it is still 3 fainter than the measured value (see Table 2). Even if we add the contribution of the upper limit of the potential broad component of H (2 3 1018 erg s-1 cm-2, Álvarez-Márquez et al. 2025), the flux in F770W would increase only by 18 27 nJy and would be still incompatible with the MIRI data by almost 2 (see Table 2). Despite the increase in the modelled flux in F1000W (i.e. 130 6 nJy) driven by the optical and near-IR emission from the older stellar population, we do not recover the observed MIRI flux (i.e. 249 77 nJy).
@@
Although assuming either a single or mixed stellar population provides a good SED fit up to rest-frame 0.5 m, both approaches fail to reproduce the F770W and F1000W photometry. As is discussed in Sect. 4.1, the significant increase in the SED of GN-z11 beyond 0.66 m cannot be reproduced solely by a combination of different stellar populations. This result suggests the presence of an additional, intrinsically red, component in GN-z11.